Steve races to the nearest taco stand at lunchtime and sees that his pedometer recorded his peak speed at 87.7 cm/s. What was Steve's peak speed in kilometers per hour?

1
Expert's answer
2022-02-06T11:23:04-0500

"87.7\frac{cm}{s}\times\frac{1\ m}{100\ cm}\times\frac{1\ km}{1000\ m}\times\frac{3600\ s}{1\ h}=3.16\frac{km}{h}"


Answer: 3.16 km/h